Fuzzy Monkey Technologies, Inc., purchased as a long-term investment $90 million of 6% bonds, dated January 1, on January 1, 2018. Management has the positive intent and ability to hold the bonds until maturity. For bonds of similar risk and maturity the market yield was 8%. The price paid for the bonds was $73 million. Interest is received semiannually on June 30 and December 31. Due to changing market conditions, the fair value of the bonds at December 31, 2018, was $80 million.

Required: 1. to 3. Prepare the relevant journal entries on the respective dates (record the interest at the effective rate). Record Fuzzy Monkey’s investment on bonds on January 1, 2018. Record the interest revenue on June 30, 2018. Record the interest revenue on December 31, 2018.

4. At what amount will Fuzzy Monkey report its investment in the December 31, 2018, balance sheet?

5. How would Fuzzy Monkey's 2018 statement of cash flows be affected by this investment?
Answer #1

Answer 1 to 3:

Date Januarv. 1, 2018 Account Title and explanation Investment in bonds Debit Credit $90,000,000 Discount on bond investment Cash $17,000,000 $73,000,000 1 (to record investment in bonds at discount Cash Discount on bond investment $2,700,000 $220,000 une. 30, 2018 2 Interest Revenue $2,920,000 (to record interest revenue, cash received and discount amortization $2,700,000 $228,800 December. 31, 2018 Cash Discount on bond investment Interest Revenue $2,928,800 (to record interest revenue, cash received and discount amortization)

Workings:

Jan. 1, 2018:

Discount = Face value - Price paid = $90,000,000 - $73,000,000 = $17,000,000

June 30, 2018:

Interest revenue (interest at the effective rate) = $73,000,000 * 8%/2 = $2,920,000

Amortization of discount = Interest revenue - Cash received = $2,920,000 - ($90,000,000 * 6%/2) =$220,000

Carrying value of bond investment = $73,000,000 + $220,000 = $73,220,000

December 31, 2018:

Interest revenue (interest at the effective rate) = $73,220,000* 8%/2 = $2,928,800

Amortization of discount = Interest revenue - Cash received = $2,928,800 - ($90,000,000 * 6%/2) =$228,800

Carrying value of bond investment = $73,220,000 + $228,800 = $73,448,800

Answer 4:

Face value = $90,000,000

Less, contra entry = Discount on bond investment = $17,000,000 - 220,000 - 228,800 =$16,551,200

Carrying value of bond investment = $90,000,000 - $16,551,200 = $73,448,800

Amount will Fuzzy Monkey report its investment in the December 31, 2018, balance sheet = $73,448,800

Answer 5:

Effect on 2018 statement of cash flows

Operating cash flow:

If direct method is used:

Cash inflow from interest revenue = $2,700,000 + $2,700,000 = $5,400,000

If indirect method is used = Cash inflow from interest revenue = Interest revenue - Amortization of discount = ($2,920,000 + $2,928,800 - $220,000 - $228,800) = $5,400,000

Investing cash flows:

Cash outflow from purchasing investments = $73,000,000